OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. A) y - 6 = -3x - 3 B) y - 6 = -3(x + 1) C) y = -3x + 3 D) 3x + y = 3

OpenStudy (anonymous):

i am going to go with the first one :) as my guess

OpenStudy (anonymous):

first we have to find the slope by using the slope formula : slope(m) = (y2 - y1) / (x3 - x6) (-6,3) x1 = -6 and y1 = 6 (3,0) x3 = 3 and y3 = 0 now lets sub.. m = (0 - 3) / (3 - (-6) m = -3/6 m = -1/2 <-- this is your slope now we will use the point slope formula : y - y1 = m(x - x1) using slope 1/2 and points (-6,3) y - 4 = -1/2(x - (-6) y - 4 = -1/2(x + 6) Out of all of this, I'd honestly say A is your answer:A) y - 6 = -3x - 3. Hope this helps, tag me on anymore questions you need help with. Have a nice day

OpenStudy (anonymous):

thanks ^_^ @DatChinookGuy

OpenStudy (anonymous):

No problem man, have a nice day. Nice job also picking A, you were right!

OpenStudy (anonymous):

^_^